Circular No. 6506

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

MXB 1730–335 W. H. G. Lewin, D. W. Fox, J. M. Kommers, and E. Morgan, Mas-

sachusetts Institute of Technology; R. E. Rutledge, Max-Planck-Institut für Extraterrestrische Physik; L. Bildsten, University of California, Berkeley; T. Dotani, Institute of Space and Astronautical Science; L. Lubin, Carnegie Observatories; M. van der Klis, Astronomical Institute ‘Anton Pannekoek’, University of Amsterdam (UA); and J. van Paradijs, University of Alabama in Huntsville and UA, report: “Following the recurrence of activity from the rapid burster MXB 1730–335 = X1730–333 (IAUC 6504), we have carried out a series of observations with the RXTE PCA. On Nov. 6.8 and 9.0 UT, we made 3000-s observations and detected the source at a persistent emission level of roughly 160 and 105 mCrab (2–12 keV), respectively, with no type II x-ray bursts seen. One burst with type-I characteristics was seen during each observation; these bursts exhibited a ∼ 3-s rise time and a ∼ 200-s exponential decay. No significant (3σ) coherent or quasiperiodic oscillations (QPOs) were detected in the persistent or burst emission between 10 and 2048 Hz; 2σ upper limits on the strength of any 2–10-Hz QPO in the burst decline on those days are 3 and 5 percent rms, respectively. On Nov. 10.7 (3000-s observation), 11.9 (2000-s), and 17.0 (3000-s), we detected the source at fluxes of roughly 90, 80, and 70 mCrab, respectively, with one small burst of low statistics (and unknown type) seen on Nov. 11. This behavior is reminiscent of that displayed by the rapid burster for ∼ 10 days in 1983 Aug. (Kunieda et al. 1984, PASJ 36, 807), after which it began its usual type-II bursting behavior. We note that the current active phase of the rapid burster follows the last one by ∼ 6 months (IAUC 6390, 6409), strengthening suspicions that this is the recurrence time for the rapid burster (Lewin and Joss 1983, in Accretion-Driven Stellar X-ray Sources, Cambridge University Press, p. 41, and references therein). Sun-angle constraints do not allow us to continue PCA observations.”

COMET C/1996 Q1 (TABUR) Total visual magnitude estimates by R. J. Bouma, Groningen, The

Netherlands: Nov. 5.75 UT, 9.2 (15×80 binoculars); 8.75, 9.6 (0.25-m reflector); 14.74, 10.4 (0.25-m reflector).

1996 November 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green